# SAC Flow: Sample-Efficient Reinforcement Learning of Flow-Based Policies via Velocity-Reparameterized Sequential Modeling

### v1

@@ Abstract
Training expressive flow-based policies with off-policy reinforcement learning is notoriously unstable due to gradient pathologies in the multi-step action sampling process. We trace this instability to a fundamental connection: the flow rollout is algebraically equivalent to a residual recurrent computation, making it susceptible to the same vanishing and exploding gradients as RNNs. To address this, we reparameterize the velocity network using principles from modern sequential models, introducing two stable architectures: Flow-G, which incorporates a gated velocity, and Flow-T, which utilizes a decoded velocity. We then develop a practical SAC-based algorithm, enabled by a noise-augmented rollout, that facilitates direct end-to-end training of these policies. Our approach supports both from-scratch and offline-to-online learning and achieves state-of-the-art performance on continuous control and robotic manipulation benchmarks, eliminating the need for common workarounds like policy distillation or surrogate objectives. Anonymized code is available at https://github.com/Elessar123/SAC-FLOW.
@@ 1 Introduction
  • State-of-the-art sample efficiency and performance. Our proposed methods, SAC Flow-G and SAC Flow-T, consistently attain state-of-the-art performance across a comprehensive suite of benchmarks, including MuJoCo, OGBench, and Robomimic. They significantly outperform recent flow- and diffusion-based baselines in terms of both final returns and convergence speed, empirically validating the high sample efficiency of our direct off-policy training approach. Ablation studies further confirm the robustness and stability of our designs.
@@ 4 Training Flow-based Policy via SAC
For clarity, we refer to our methods as SAC Flow-G and SAC Flow-T, corresponding to training with Flow-G and Flow-T via SAC, respectively. Both terms apply to both from-scratch and offline-to-online training variants.
@@ 5.1.1 Environments and Offline Datasets
Figure 4: From-scratch training performance. Our SAC Flow-T and SAC Flow-G consistently outperform all baselines across MuJoCo tasks (Fig. (a)-(f)), achieving significant sample efficiency and convergence stability. However, all methods struggle on the hard-exploration, sparse-reward tasks (Can from Robomimic, and Cube-Double from OGBench), highlighting the necessity of offline-to-online training.
@@ 5.2 Main results
Fig. 4 presents the results for from-scratch training. Our methods, SAC Flow-G and SAC Flow-T, achieve the state-of-the-art performance in all MuJoCo tasks. We observe that all flow-based baselines generally converge faster than diffusion-based baselines, where SAC Flow-T, SAC Flow-G and FlowRL perform better than DIME and QSM. Furthermore, benefiting from direct optimization with the SAC objective, our methods also surpass FlowRL, which relies on a surrogate objective with a Wasserstein distance constraint. Meanwhile, with the expressive parameterization of flow-based policy, our method achieves much higher final performance in challenging tasks (Fig. 4(d), 4(e), and 4(f)), and remains convergence stability in simple tasks (Fig. 4(a), 4(b), and 4(c)). For reference, we include the on-policy baseline, PPO, to highlight the superior sample efficiency of off-policy algorithms. However, from-scratch training struggles in tasks with large exploration spaces and sparse rewards (Figs. 4(g), 4(h)), underscoring the necessity of an offline-to-online training setting.
Fig. 5 shows the offline-to-online training performance in sparse reward tasks. All methods are trained on 1M offline updates followed by 1M online steps. In the challenging OGBench environments, including cube-triple and cube-quadruple, our proposed methods, particularly SAC Flow-T, achieve rapid convergence and attain a state-of-the-art overall success rate. In the Robomimic environment, however, SAC Flow-T and SAC Flow-G only yield results comparable to QC-FQL. This is primarily because the training is strictly regularized with a large $\beta$ value (Equ. (12)). As a result, the learning capacity of the flow model is severely limited, causing its performance to be similar to that of the one-step policy in QC-FQL. We further compare the on-policy baseline, Reinflow, in Robomimic. Leveraging the high data efficiency of off-policy learning, our SAC Flow-G and SAC Flow-T outperform Reinflow under 1M online steps. The additional results are available in Appendix E.

### v2

@@ Abstract
Training expressive flow-based policies with off-policy reinforcement learning is notoriously unstable due to gradient pathologies in the multi-step action sampling process. We trace this instability to a fundamental connection: the flow rollout is algebraically equivalent to a residual recurrent computation, making it susceptible to the same vanishing and exploding gradients as RNNs. To address this, we reparameterize the velocity network using principles from modern sequential models, introducing two stable architectures: Flow-G, which incorporates a gated velocity, and Flow-T, which utilizes a decoded velocity. We then develop a practical SAC-based algorithm, enabled by a noise-augmented rollout, that facilitates direct end-to-end training of these policies. Our approach supports both from-scratch and offline-to-online learning and achieves state-of-the-art performance on continuous control and robotic manipulation benchmarks, eliminating the need for common workarounds like policy distillation or surrogate objectives. Code is available at https://github.com/Elessar123/SAC-FLOW.
@@ 1 Introduction
  • State-of-the-art sample efficiency and performance. Our proposed methods, SAC Flow-G and SAC Flow-T, establish new state-of-the-art results across a comprehensive suite of benchmarks. In from-scratch training on challenging MuJoCo tasks, our approach delivers performance gains of up to $130\%$ over strong baselines. Furthermore, in complex offline-to-online manipulation tasks on OGBench, it achieves up to a $60\%$ higher success rate. These results empirically validate the superior sample efficiency of our direct off-policy training approach, with ablation studies further confirming the robustness of our designs.
@@ 4 Training Flow-based Policy via SAC
For clarity, we refer to our methods as SAC Flow-G and SAC Flow-T, corresponding to training our Flow-G and Flow-T architectures with SAC. These terms apply to both our from-scratch and offline-to-online training procedures. Additionally, we extend our framework to fine-tuning, where Flow-G and Flow-T are reformulated as lightweight adapters to stabilize the online fine-tuning of an arbitrary pre-trained flow-based policy. This adapter-based fine-tuning algorithm is detailed in Appendix F.
@@ 5.1.1 Environments and Offline Datasets
Figure 4: From-scratch training performance. Our SAC Flow-T and SAC Flow-G achieve comparable or better performance accross all tasks except Humanoid (Fig. (a)-(f)), demonstrating significant sample efficiency and convergence stability. However, all methods struggle on the hard-exploration, sparse-reward tasks (Can from Robomimic, and Cube-Double from OGBench), highlighting the necessity of offline-to-online training.
@@ 5.2 Main results
Figure 4 illustrates the results for from-scratch training. Our methods, SAC Flow-G and SAC Flow-T, achieve superior or comparable performance across most MuJoCo tasks, with the exception of Humanoid. Although DIME and FlowRL generally converge faster than other baselines, our methods consistently surpass FlowRL, benefiting from direct optimization of the SAC objective. Furthermore, SAC Flow outperforms DIME in Hopper (Fig. 4(a)), Walker (Fig. 4(b)), and HumanoidStandup (Fig. 4(f)), while achieving comparable results in HalfCheetah (Fig. 4(c)) and Ant (Fig. 4(d)). Moreover, with the expressive parameterization of flow-based policy, our method achieves much higher final performance in challenging tasks, demonstrating up to a $130\%$ improvement over the baseline ( 4(f)), and remains convergence stability in simple tasks (Fig. 4(a), 4(b), and 4(c)). For reference, we include the on-policy baseline, PPO, to highlight the superior sample efficiency of off-policy algorithms. However, all from-scratch methods struggle in tasks with large exploration spaces and sparse rewards (Figs. 4(g), 4(h)), underscoring the necessity of an offline-to-online training setting.
Fig. 5 shows the offline-to-online training performance in sparse reward tasks. All methods are trained on 1M offline updates followed by 1M online steps. In the challenging OGBench environments, including cube-triple and cube-quadruple, our proposed methods, particularly SAC Flow-T, achieve rapid convergence and attain a state-of-the-art overall success rate, demonstrating up to a $60\%$ improvement over the baseline. In the Robomimic environment, however, SAC Flow-T and SAC Flow-G only yield results comparable to QC-FQL. This is primarily because the training is strictly regularized with a large $\beta$ value (Equ. (12)). As a result, the learning capacity of the flow model is severely limited, causing its performance to be similar to that of the one-step policy in QC-FQL. We further compare the on-policy baseline, Reinflow, in Robomimic. Leveraging the high data efficiency of off-policy learning, our SAC Flow-G and SAC Flow-T outperform Reinflow under 1M online steps. The additional results are available in Appendix E.
@@ Appendix E More Experiment Results
Appendix F The Algorithm Framework for Fine-tuning Setting
In this section, we provide a detailed description of how our Flow-G and Flow-T frameworks can be adapted to fine-tune an arbitrary pre-trained flow-based policy, denoted $v_{\theta}^{pre}$. We introduce specific formulations and initialization strategies designed to seamlessly integrate $v_{\theta}^{pre}$ while preserving its initial performance.
Flow-G (Adapter): We utilize a simplified variant of the gated velocity from Equ. (6). Instead of $v_{\theta}=g_{i}\odot\left(\hat{v}_{\theta}-A_{t_{i}}\right)$, we define the velocity as $v_{\theta}=g_{i}\odot\hat{v}_{\theta}$, where $g_{i}={Sig}\,\big(z_{\theta}(t_{i},A_{t_{i}},s)\big)$ acts as a modulating gate. We set the candidate network $\hat{v}_{\theta}$ to be the frozen, pre-trained network $v_{\theta}^{pre}$. The new gate network $z_{\theta}$ is initialized such that its output bias is large (e.g., 5.0) and weights are zero. At the beginning of fine-tuning, this causes $g_{i}\approx 1$, making the gate act as an identity mapping ($v_{\theta}\approx v_{\theta}^{pre}$). The fine-tuning process then only optimizes the lightweight gate network $z_{\theta}$.
Flow-T (Adapter): For Flow-T, we adapt the decoder architecture from Equ. (8). We replace the final linear projection $W_{o}$ with the entire pre-trained network $v_{\theta}^{pre}$. The reformulated velocity becomes $v_{\theta}(t_{i},A_{t_{i}},s)=v_{\theta}^{pre}\Big(\mathrm{LN}(\Phi_{A_{i}}^{(L)})\Big)$. The newly introduced cross-attention and feed-forward blocks within the Transformer stack are zero-initialized. Due to the residual connections, the output $\mathrm{LN}(\Phi_{A_{i}}^{(L)})$ initially passes through the unchanged action-time embedding, effectively preserving the original input structure expected by $v_{\theta}^{pre}$. This ensures $v_{\theta}$ behaves identically to $v_{\theta}^{pre}$ at the start of fine-tuning.
In most fine-tuning scenarios, the critic $Q_{\psi}$ is initialized from scratch. To allow the critic to learn meaningful values before the policy is updated with the RL objective, and to let the adapter networks settle, we introduce a ”warm-up” phase. During this phase, the critic is trained normally via Equ. (11), but the actor is trained only with a behavioral cloning loss:
$L_{\text{actor}}^{warm}(\theta)=\beta\,\|a_{h}^{\theta}-a_{h}\|_{2}^{2},\quad(s_{h},a_{h})\sim\mathcal{B},\quad a_{h}^{\theta}=\tanh(A_{t_{K}}^{\theta})\sim\pi_{\theta}(\cdot\mid s_{h}).$ (37)
Using these specific adapter formulations, we introduce the fine-tuning algorithm in Algo. 3. The reformulated policy is first trained for $L_{\text{warmup}}$ steps, followed by $L_{\text{online}}$ steps of full online training. The experimental results are shown in Fig. 16 and Fig. 17. The fine-tuning methods achieve performance comparable to the offline-to-online methods. It is surprising that the fine-tuning methods achieve better performance in Cube-Triple-Task3, Task4, and Robomimic-Square. This suggests that a randomly initialized critic might be easier to optimize during the online phase than one pre-trained via offline RL.
Algorithm 3 SAC Flow (fine-tuning)
1: Initialize critic $Q_{\psi}$, target $Q_{\bar{\psi}}$. Set $\mathcal{B}\leftarrow\mathcal{D}_{\text{expert}}$.
2: Load pre-trained $v_{\theta}^{pre}$ and construct $\pi_{\theta}$ using the Flow-G or Flow-T adapter formulation.
3:for $\ell=1$ to $L_{\text{warmup}}$ do
5:  Sample $\{(s_{h},a_{h},r_{h},s_{h+1})\}_{h=1}^{N}\sim\mathcal{B}$.
6:  Actor: minimize Equ. (37) with $a_{h}^{\theta}$ from the noisy rollout.
7:  Critic: minimize Equ. (11); update the target network.
8:end for
9:for $\ell=L_{\text{warmup}}+1$ to $L_{\text{warmup}}+L_{\text{online}}$ do
10:  Interact with the environment using $\pi_{\theta}$; append to $\mathcal{B}$.
11:  Sample $\{(s_{h},a_{h},r_{h},s_{h+1})\}_{h=1}^{N}\sim\mathcal{B}$.
12:  Actor: minimize Equ. (12) with $a_{h}^{\theta}$ from the noisy rollout.
13:  Critic: minimize Equ. (11); update the target network.
14:end for
Figure 16: Complete fine-tuning performance on OGBench. All methods are pre-trained for 1M steps using flow matching with a standard policy. The pre-trained policy is then adapted using our Flow-G or Flow-T formulation and fine-tuned, starting with 0.1M warm-up steps followed by 1M online interaction steps (see Algo. 3).
Figure 17: Complete fine-tuning performance on Robomimic. The training protocol is identical to the OGBench experiments (Fig. 16): 1M steps of flow matching pre-training, followed by 0.1M warm-up steps and 1M online fine-tuning steps.